OpenStudy (anonymous):

Find all solutions in the interval [0, 2π) cos2x + 2 cos x + 1 = 0

OpenStudy (kainui):

Wait, isn't this the same question as the last one, except it's cos instead of sin?

OpenStudy (anonymous):

no i don't think so

OpenStudy (kainui):

Yeah, you can definitely factor this. that's a cos^2(x) right?

OpenStudy (anonymous):

yea

OpenStudy (kainui):

remember cos^2(x)=(cosx)^2. Just plug in something like y=cosx into the formula and factor it out like normal.

OpenStudy (anonymous):

i got x=2pi

OpenStudy (kainui):

Show me your steps.

OpenStudy (anonymous):

my bad i got x= pi/4, 7pi/4

OpenStudy (anonymous):

wrong problem

OpenStudy (abb0t):

\[\cos(2x) = 2\cos^2(x)-1\] Now, you have \[2\cos^2(x)-1 + 2 \cos(x) + 1 = 0\] Which is similar to: \[ax^2 + bx + c = 0\]

OpenStudy (kainui):

abb0t did you read, he said it was cos^2(x) not cos(2x)

OpenStudy (abb0t):

Didn't read.

OpenStudy (anonymous):

so... i tried it again and i got x= pi/2, 3pi/2

OpenStudy (kainui):

Why don't you write out your steps and I can help you along where you need it, if you do.

OpenStudy (anonymous):

x=pi i got it

OpenStudy (kainui):

Well I'm not going to offer you any help if you're not going to try the problem.